# Unraveling-induced entanglement phase transition in diffusive trajectories of continuously monitored noninteracting fermionic systems

### v1

@@ Abstract
Introduction.— Entanglement stands out as the most paradigmatic feature of quantum mechanics. Beyond its relevance in quantum information, the spreading of entanglement in many-body systems is tied to fundamental questions [1], e.g., related to the emergence of critical correlations close to quantum phase transitions [2, 3, 4, 5, 6] or of thermal ensembles in isolated quantum systems [7, 8, 9, 10]. To shed light on these phenomena, basic models of random quantum circuits have been introduced [11, 12]. Their analysis demonstrated that generic (nonintegrable) unitary systems evolve towards strongly correlated states, displaying volume-law entanglement [13]. That is, bipartite entanglement which grows with the size $\ell$ of the smallest subsystem generated by a bipartition [see sketch in Fig. 1(a)].
@@
Observing local properties of many-body quantum systems in real time, either through projective measurements [11, 18, 19, 20, 21, 22, 23] or through (weak) continuous-monitoring processes, can dramatically affect the built-up of entanglement [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. For a small measurement rate $\gamma$, the unitary volume-law behavior may be expected to survive the presence of a local monitoring. However, for a large rate $\gamma$, the quantum state necessarily remains close to a product state and thus features area-law entanglement, scaling with the size of the boundary of a bipartition [cf. Fig. 1(a)]. Remarkably, many-body systems can display a genuine transition between these two phases, as a function of the measurement rate $\gamma$ [30, 31, 32, 32, 38, 33, 39, 40]. For the case of noninteracting fermionic systems, measurement-induced transitions can occur from an area-law phase to one featuring entanglement growth that is logarithmic in $\ell$ [26, 41].
In this work, we take a different perspective on entanglement phase transitions. Rather than exploring the behavior of the system upon varying the measurement rate $\gamma$, we analyze the entanglement phases generated by varying the measurement process itself [see sketch in Fig. 1(a)]. The rationale is as follows. The dynamics of the quantum state averaged over all realizations of a continuous-monitoring process is described by a Markovian quantum master equation. Different monitoring processes, even when performed at the same rate $\gamma$, are associated with different “unravelings” of the quantum master equation into quantum stochastic processes. It is thus important to understand how entanglement properties of many-body systems depend on the considered measurement process, or unraveling. In Refs. [36, 37], this dependence was explored in the context of devising optimal simulation strategies, where an entanglement transition driven by a change of the unraveling dynamics was observed. However, the system considered there is a fully random quantum circuit, for which the unraveling transition could be exactly mapped onto a measurement-induced one [36, 37]. The question thus remains whether an unraveling-induced phase transitions can occur for many-body systems possessing a deterministic Hamiltonian evolution and for which a mapping to a standard measurement-induced transition is not possible.
@@
Monitored tight-binding model.— Our study is based on a one-dimensional tight-binding lattice model, with periodic boundary conditions, which is a paradigmatic system for the study of transport and correlation spreading. The system, see Fig. 1(a), is made of $L$ sites, each one hosting a fermionic particle with the corresponding creation and annihilation operators $c_{i}^{\dagger},c_{i}$. The Hamiltonian dynamics entails nearest-neighbor hopping represented by the operator
@@
which provides the increment of the system state, ${\rm d}\ket{\psi_{t}}$, during an infinitesimal time interval ${\rm d}t$. The term ${\rm d}W_{t}^{i}$ is the increment of a standard Wiener process, obeying $\mathbb{E}[{\rm d}W_{t}^{i}{\rm d}W_{t}^{j}]=\delta_{ij}{\rm d}t$ and $\mathbb{E}[{\rm d}W_{t}^{i}]=0$, with $\mathbb{E}$ denoting expectation over the processes. We have further defined the operator $\Delta_{t}n_{i}=\left[n_{i}-\langle n_{i}\rangle_{t}\right]$. The process in Eq. (4) is used to model, for instance, homodyne-detection measurements [42], with $\gamma_{\rm I}$ encoding the monitoring rate. For a large ratio $\gamma_{\rm I}/\lambda$, the average stationary entanglement entropy $\bar{S}_{\infty}$ of the process saturates to a value which does not depend on system size, indicative of an area-law behavior as shown in Fig. 2(a) [26, 28]. Conversely, for small values $\gamma_{\rm I}/\lambda$, the system enters a phase with subextensive stationary entanglement entropy, growing as the logarithm of the system size [26, 27]. The unraveling in Eq. (4), therefore, does not feature volume-law entanglement [26, 28].
@@
with ${\rm d}V_{t}^{i}$ being a standard Wiener increment. Such a process is also related to homodyne detection [42] and $\gamma_{\rm II}$ is the monitoring rate. Due to the imaginary unit in front of the noise terms, the process in Eq. (5) is generated by a stochastic Hamiltonian and has effectively no nontrivial non-Hermitean component. Essentially, Eq. (5) describes the system dynamics in the presence of a fluctuating on-site potential described by independent Wiener increments. As clearly displayed in Fig. 2(b), in this case bipartite entanglement grows as $\bar{S}_{t}\sim\sqrt{\lambda t}$ and approaches a saturation value $\bar{S}_{\infty}$ displaying volume-law behavior.
Interpolating between unravelings.— Evidently, the stochastic processes in Eq. (4) and in Eq. (5) result in qualitatively different scaling of entanglement. In the following, we want to explore whether interpolating between the two, while keeping the measurement rate fixed, can result in an entanglement phase transition. By examining Fig. 2(a) and Fig. 2(b), one may expect a transition from area-law to volume-law behavior. However, as we will show, at most logarithmic entanglement scaling is possible.
@@
Before presenting our results for the different stochastic processes generated by Eq. (6), we discuss the numerical procedure we exploit [28, 26, 43]. Since the dynamics in Eq. (6) is implemented by a quadratic operator which conserves the total number of particles, we can write the state of the system, at any time $t$, as
@@
Here, $N$ is the number of fermionic particles (in our case $N=L/2$) and $U$ is an $L\times N$ isometry, $U^{\dagger}U=\mathds{1}_{N}$ where $\mathds{1}_{N}$ is the identity matrix of dimension $N$. To analyze the dynamics of the quantum state, it is sufficient to understand how the matrix $U(t)$ evolves. By considering Eq. (6) and by discretizing time, we can find the operator $K_{t}$ such that |$\ket{\psi_{t+{\rm d}t}}\approx e^{K_{t}}\ket{\psi_{t}}$. One can check that $K_{t}\ket{0}=0$ and that $K_{t}$ is quadratic. As such, the evolution of $U(t)$ can be obtained by calculating the linear combinations of creation operators generated by $e^{K_{t}}c_{j}^{\dagger}e^{-K_{t}}=\sum_{m=1}^{L}R_{jm}c_{m}^{\dagger}$. This provides an update rule for the matrix $U(t)\to U(t+{\rm d}t)$ and we set $\gamma{\rm d}t=0.1$ in our numerical simulations. The isometry $U(t)$ gives access to the fermionic single-particle correlation function $D_{mn}(t)=\langle c^{\dagger}_{m}c_{n}\rangle_{t}=\left[U(t)U^{\dagger}(t)\right]^{*}_{mn}$, from which we can calculate the entanglement entropy [44, 45, 28, 26] (see details in the Supplemental Material [46]). We average this quantity over several realizations of the stochastic processes and estimate its stationary behavior by integrating over a finite window at large times. By construction, the entanglement entropy behaves as shown in Fig. 2(a,b), for the limiting cases $\alpha=0$ and $\alpha=1$, showing area-law and volume-law entanglement, respectively.
@@
Figure 3: Unraveling-induced entanglement phase transition. (a) Stationary values of the entanglement entropy $\bar{S}_{\infty}(\alpha,L)$ as a function of the logarithm of the system size, for different values of $\alpha$. The inset shows the data in linear scale. For small parameter $\alpha$, $\bar{S}_{\infty}(\alpha,L)$ exhibits an area-law behavior. Upon increasing $\alpha$, $\bar{S}_{\infty}(\alpha,L)$ starts to display a subextensive logarithmic growth with the system size. (b) Estimate of the order parameter $c(\alpha)$ [see Eq. (8)], obtained as $\bar{S}_{\infty}(\alpha,L)/\log L$, for $L\in\left\{120,250,400,600,700,800\right\}$. The dotted line represents $c(\alpha)$ from a fit of Eq. (8) using the largest four system sizes considered in panel (a). The inset shows instead the fit of the residual entropy $s_{0}(\alpha)$. We have considered $\lambda=\gamma/2$ and the average is performed over $100$ realizations of the process.
@@
Unraveling phase transition.— In the regime $10<\alpha<$ 1, we see that no volume-law entanglement scaling is possible [cf. Fig. 3(a)]. Nonetheless, two different regimes emerge. For $\alpha$ smaller than a critical value $\alpha_{\rm c}$, the stochastic process features an area-law entanglement behavior, $\bar{S}_{\infty}\propto O(1)$ as in the case of $\alpha=0$. On the other hand, for $\alpha>\alpha_{\rm c}$ there appears a subextensive logarithmic correction, in the system size, such that $\bar{S}_{\infty}\propto\log L$. The latter is reminiscent of the one observed in Ref. [26] upon variations of the measurement rate.
@@
The presence of an extended logarithmic regime, reminiscent of a conformally invariant critical phase in purely Hamiltonian dynamics [47, 6, 48], is suggestive of Berenzinskii-Kosterlitz-Thouless (BKT) universal behavior [49] close to the transition point [26]. Under this assumption, we can estimate the critical unraveling parameter which separates the two entanglement regimes. First of all, the value of $\alpha$ at which the $s_{0}(\alpha)$ changes sign, from positive to negative, may provide an estimate for the critical unraveling parameter $\alpha_{\rm c}$ [26]. To consolidate this estimate, which from the inset of Fig. 3(b) appears to be roughly $\alpha_{\rm c}\approx 0.8$, we explore a finite-size scaling analysis for $\bar{S}_{\infty}(\alpha,L)$ [26, 50, 51]. We consider the function $\bar{S}_{\infty}\left(\alpha,{L}\right)-\bar{S}_{\infty}\left(\alpha_{c},{L}\right)$ and plot it as a function of the quantity $\left(\alpha-\alpha_{c}\right)\log^{2}L$ for different values of $\alpha_{\rm c}$. When considering the exact value of the critical parameter, one would expect to observe a collapse of all data points onto a unique scaling function. In Fig. 4 we report the scaling that we have obtained by assuming $\alpha_{\rm c}=0.8$. For the latter value, the finite-size scaling works reasonably well, given also unavoidable finite-size effects, suggesting that the transition is within the BKT universality. The inset of Fig. 4 shows the bare data for $\bar{S}_{\infty}\left(\alpha,L\right)$ as a function of $\alpha$.
@@
Conclusions.— We have considered a family of continuous-monitoring processes identified by an unraveling parameter $\alpha$. We have shown that varying the measurement process (achieved by varying $\alpha$) can have a dramatic impact on the spreading of entanglement in the considered noninteracting tight-binding model. When the unraveling parameter reaches a critical value, the system transitions from a phase characterized by an area-law to one exhibiting logarithmic entanglement behavior. The latter phase is reminiscent of emergent conformal invariance in unitary systems and the entanglement phase transition indeed appears to belong to the BKT universality [26]. It would be interesting to study the entanglement phase diagram in the $\gamma-\alpha$ plane [36]. We expect that for sufficiently small $\gamma$ the system will always be found in the logarithmic phase. On the other hand, we expect that for large $\gamma$, the value of $\alpha_{\rm c}$, determining the onset of the logarithmic phase, moves towards $\alpha_{\rm c}\to 1$.
@@
In Ref. [36] an unraveling transition was observed in a fully random quantum circuit, where one of the two “extremal” unravelings could be absorbed in the random Hamiltonian ensemble. In contrast, our system features a deterministic Hamiltonian evolution and two unravelings both substantially altering the unitary dynamics. As a result, the entanglement phase transition we observe cannot be directly mapped onto a “standard” measurement-induced one [36]. Signatures of our unraveling-induced entanglement phase transition may be observed in experiments with cold-atoms [42], for instance by investigating the behavior of many-body correlations [26]. Experimentally, different unravelings may be obtained by varying the quadrature component of the monitored output field [42]. This possibility may allow for controlling entanglement in many-body states via continuous monitoring, which is potentially relevant for tailoring quantum correlations in noisy devices.

### v2

@@ Abstract
I Introduction
Entanglement stands out as the most paradigmatic feature of quantum mechanics. Beyond its relevance in quantum information, the spreading of entanglement in many-body systems is tied to fundamental questions [1], e.g., related to the emergence of critical correlations close to quantum phase transitions [2, 3, 4, 5] or of thermal ensembles in isolated quantum systems [6, 7, 8]. To shed light on these phenomena, basic models of random quantum circuits have been introduced [9, 10, 11, 12, 13, 14]. Their analysis demonstrated that generic (nonintegrable) unitary systems evolve towards strongly correlated states, displaying volume-law entanglement [15]. That is, bipartite entanglement which grows with the size $\ell$ of the smallest subsystem generated by a bipartition [see sketch in Fig. 1(a)].
Observing local properties of many-body quantum systems in real time, either through projective measurements [9, 16, 17, 18, 19, 20, 21, 22] or through (weak) continuous-monitoring processes [23], can dramatically affect the built-up of entanglement[17, 16, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. For a small measurement rate $\gamma$, the unitary volume-law behavior may be expected to survive the presence of a local monitoring. However, for a large rate $\gamma$, the quantum state necessarily remains close to a product state and thus features area-law entanglement, scaling with the size of the boundary of a bipartition [cf. Fig. 1(a)]. Remarkably, many-body systems can display a genuine transition between these two phases, as a function of the measurement rate $\gamma$ [26, 36, 36, 39, 24, 40, 41, 38]. For the case of noninteracting fermionic systems, measurement-induced transitions can occur from an area-law phase to one featuring entanglement growth that is logarithmic in $\ell$ [27, 42].
In this work, we take a different perspective on entanglement phase transitions. Rather than exploring the behavior of the system upon varying the measurement rate $\gamma$, we analyze the entanglement phases generated by varying the measurement process itself [see sketch in Fig. 1(a)]. The rationale is as follows. The dynamics of the quantum state averaged over all realizations of a continuous-monitoring process is described by a Markovian quantum master equation. Different monitoring processes, even when performed at the same rate $\gamma$, are associated with different “unravelings” of the quantum master equation into quantum stochastic processes. It is thus important to understand how entanglement properties of many-body systems depend on the considered measurement process, or unraveling. In Refs. [32, 37], this dependence was explored in the context of devising optimal simulation strategies, where an entanglement transition driven by a change of the unraveling dynamics was observed. However, the system considered there is a quantum circuit with a fully random Hamiltonian, for which one of the two unravelings of the measurement process can be re-absorbed in the Hamiltonian ensemble without modifying it. As such, the emergent transition can be exactly mapped on a measurement-induced one since there is effectively a single relevant unravelling of the dissipative processes [32, 37]. The question thus remains whether an unraveling-induced phase transitions can occur for many-body systems possessing a deterministic Hamiltonian evolution and for which a mapping to a standard measurement-induced transition is not possible.
@@ I Introduction
II Monitored tight-binding model
Our study is based on a one-dimensional tight-binding lattice model, with periodic boundary conditions, which is a paradigmatic system for the study of transport and correlation spreading. The system, see Fig. 1(a), is made of $L$ sites, each one hosting a fermionic particle with the corresponding creation and annihilation operators $c_{i}^{\dagger},c_{i}$. The Hamiltonian dynamics entails nearest-neighbor hopping represented by the operator
@@ II Monitored tight-binding model
which provides the increment of the system state, ${\rm d}\ket{\psi_{t}}$, during an infinitesimal time interval ${\rm d}t$. The term ${\rm d}W_{t}^{i}$ is the increment of a standard Wiener process, obeying $\mathbb{E}[{\rm d}W_{t}^{i}{\rm d}W_{t}^{j}]=\delta_{ij}{\rm d}t$ and $\mathbb{E}[{\rm d}W_{t}^{i}]=0$, with $\mathbb{E}$ denoting expectation over the processes. We have further defined the operator $\Delta_{t}n_{i}=\left[n_{i}-\langle n_{i}\rangle_{t}\right]$. The process in Eq. (4) is used to model, for instance, homodyne-detection measurements [47], with $\gamma_{\rm I}$ encoding the monitoring rate. For a large ratio $\gamma_{\rm I}/\lambda$, the average entanglement entropy $\bar{S}_{\infty}$ of the process saturates to a value which does not depend on system size, indicative of an area-law behavior as shown in Fig. 2(a) [27, 25]. Conversely, for small values $\gamma_{\rm I}/\lambda$, the system enters a phase with subextensive stationary entanglement entropy, growing as the logarithm of the system size [27, 30]. The unraveling in Eq. (4), therefore, does not feature volume-law entanglement [27, 25].
@@
with ${\rm d}V_{t}^{i}$ being a standard Wiener increment. Such a process is also related to homodyne detection [47] and $\gamma_{\rm II}$ is the monitoring rate. Due to the imaginary unit in front of the noise terms, the process in Eq. (5) is generated by a stochastic Hamiltonian and has effectively no nontrivial non-Hermitean component. Essentially, Eq. (5) describes the system dynamics in the presence of a fluctuating on-site potential described by independent Wiener increments. As clearly displayed in Fig. 2(b), in this case bipartite entanglement grows as $\bar{S}_{t}\sim\sqrt{\lambda t}$ [48, 25] and approaches a saturation value $\bar{S}_{\infty}$ displaying, on average, the same volume-law behavior shown by the deterministic unitary dynamics $\gamma=0$ (see Fig. B2 in Appendix B for a comparison).
III Interpolating between unravelings
Evidently, the stochastic processes in Eq. (4) and in Eq. (5) result in qualitatively different scaling of entanglement. In the following, we want to explore whether interpolating between the two, while keeping the measurement rate fixed, can result in an entanglement phase transition. By examining Fig. 2(a) and Fig. 2(b), one may expect a transition from area-law to volume-law behavior. However, as we will show, at most logarithmic entanglement scaling is possible.
@@ III Interpolating between unravelings
Before presenting our results for the different stochastic processes generated by Eq. (6), we briefly discuss the numerical procedure we exploit [25, 27, 49] (for a detailed discussion see Appendix A). Since the dynamics in Eq. (6) is implemented by a quadratic operator which conserves the total number of particles, we can write the state of the system, at any time $t$, as
@@
Here, $N$ is the number of fermionic particles (in our case $N=L/2$) and $U$ is an $L\times N$ isometry, $U^{\dagger}U=\mathds{1}_{N}$ where $\mathds{1}_{N}$ is the identity matrix of dimension $N$. To analyze the dynamics of the quantum state, it is sufficient to understand how the matrix $U(t)$ evolves. By considering Eq. (6) and by discretizing time, we can find the operator $K_{t}$ such that |$\ket{\psi_{t+{\rm d}t}}\approx e^{K_{t}}\ket{\psi_{t}}$. One can check that $K_{t}\ket{0}=0$ and that $K_{t}$ is quadratic. As such, the evolution of $U(t)$ can be obtained by calculating the linear combinations of creation operators generated by $e^{K_{t}}c_{j}^{\dagger}e^{-K_{t}}=\sum_{m=1}^{L}R_{jm}c_{m}^{\dagger}$. This provides an update rule for the matrix $U(t)\to U(t+{\rm d}t)$ and we set $\gamma{\rm d}t=0.1$ in our numerical simulations (see also Fig. B1 in Appendix B for an analysis with $\gamma{\rm d}t=0.05$). The isometry $U(t)$ gives access to the fermionic single-particle correlation function $D_{mn}(t)=\langle c^{\dagger}_{m}c_{n}\rangle_{t}=\left[U(t)U^{\dagger}(t)\right]^{*}_{mn}$, from which we can calculate the entanglement entropy [50, 51, 25, 27] (see details in Appendix A). We average this quantity over several realizations of the stochastic processes and estimate its stationary behavior by integrating over a finite window at large times. By construction, the entanglement entropy behaves as shown in Fig. 2(a,b), for the limiting cases $\alpha=0$ and $\alpha=1$, showing area-law and volume-law entanglement, respectively.
@@
Figure 3: Unraveling-induced entanglement phase transition. (a) Stationary values of the entanglement entropy $\bar{S}_{\infty}(\alpha,L)$ as a function of the logarithm of the system size, for different values of $\alpha$. The inset shows the data in linear scale. For small parameter $\alpha$, $\bar{S}_{\infty}(\alpha,L)$ exhibits an area-law behavior. Upon increasing $\alpha$, $\bar{S}_{\infty}(\alpha,L)$ starts to display a subextensive logarithmic growth with the system size. (b) Estimate of the order parameter $c(\alpha)$ [see Eq. (8)], obtained as $\bar{S}_{\infty}(\alpha,L)/\log L$, for $L\in\left\{120,250,400,600,700,800\right\}$. The dashed line represents $c(\alpha)$ obtained from a fit of Eq. (8) to the data of the same system sizes (or the six largest ones). The inset shows instead the fit of the residual entropy $s_{0}(\alpha)$. We have considered $\lambda=\gamma/2$ and the average is performed over $100$ realizations of the process.
IV Unraveling phase transition
@@
In the regime $10<\alpha<$ 1, we see that no volume-law entanglement scaling is possible [cf. Fig. 3(a)]. Nonetheless, two different regimes emerge. For $\alpha$ smaller than a critical value $\alpha_{\rm c}$, the stochastic process features an area-law entanglement behavior, $\bar{S}_{\infty}\propto O(1)$ as in the case of $\alpha=0$. On the other hand, for $\alpha>\alpha_{\rm c}$ after a transient regime displaying a logarithmic entanglement growth in time $\propto\log\left(\gamma t\right)$ [see Fig.B2(a) in Appendix B], the system approaches a stationary behavior proportional to the logarithm of the subsystem size, $\bar{S}_{\infty}\propto\log L$. The latter is reminiscent of the one observed in Ref. [27] upon variations of the measurement rate.
@@
The presence of an extended logarithmic regime, reminiscent of a conformally invariant critical phase in purely Hamiltonian dynamics [52, 5, 53], is suggestive of Berenzinskii-Kosterlitz-Thouless (BKT) universal behavior [54] close to the transition point [27]. Under this assumption, we can estimate the critical unraveling parameter which separates the two entanglement regimes. First of all, the value of $\alpha$ at which the $s_{0}(\alpha)$ changes sign, from positive to negative, may provide an estimate for the critical unraveling parameter $\alpha_{\rm c}$ [27]. To consolidate this estimate, which from the inset of Fig. 3(b) appears to be roughly $\alpha_{\rm c}\approx 0.8$, we explore a finite-size scaling analysis for $\bar{S}_{\infty}(\alpha,L)$ [27, 55, 56]. We consider the function $\bar{S}_{\infty}\left(\alpha,{L}\right)-\bar{S}_{\infty}\left(\alpha_{c},{L}\right)$ and plot it as a function of the quantity $\left(\alpha-\alpha_{c}\right)\log^{2}L$ for different values of $\alpha_{\rm c}$. When considering the exact value of the critical parameter, one would expect to observe a collapse of all data points onto a unique scaling function. In Fig. 4 we report the scaling that we have obtained by assuming $\alpha_{\rm c}=0.8$. For the latter value, the finite-size scaling works reasonably well, given also unavoidable finite-size effects, suggesting that the transition is within the BKT universality. We also report a more systematic analysis of the scaling in Appendix B [cf. Fig. B3] leading to a value $\alpha_{\rm c}=0.875$. The latter, given finite-size effects and our finite data set, is in reasonable agreement with the value providing the scaling in Fig. 4. The inset of Fig. 4 shows the bare data for $\bar{S}_{\infty}\left(\alpha,L\right)$ as a function of $\alpha$.
V Conclusions
@@
We have considered a family of continuous-monitoring processes identified by an unraveling parameter $\alpha$. We have shown that varying the measurement process (achieved by varying $\alpha$) can have a dramatic impact on the spreading of entanglement in the considered noninteracting tight-binding model. When the unraveling parameter reaches a critical value, the system transitions from a phase characterized by an area-law to one exhibiting logarithmic entanglement behavior. The latter phase is reminiscent of emergent conformal invariance in unitary systems and the entanglement phase transition indeed appears to belong to the BKT universality [27]. It would be interesting to study the entanglement phase diagram in the $\gamma-\alpha$ plane [32]. We expect that for sufficiently small $\gamma$ the system will always be found in the logarithmic phase. On the other hand, we expect that for large $\gamma$, the value of $\alpha_{\rm c}$, determining the onset of the logarithmic phase, moves towards $\alpha_{\rm c}\to 1$. We note that the critical value $\alpha_{\rm c}$ that we estimate for our parameters seems to be in line with what could be obtained by considering only the (adimensional) effective measurement rate $(\gamma/\Omega)(1-\alpha)$ and comparing it with the critical value of Ref. [27]. To understand whether this is just a coincidence of whether the rate $\gamma_{\rm II}$ is really “irrelevant” for the transition, one would need indeed to derive the full phase diagram previously mentioned.
The nature and even the existence of a phase with logarithmic entanglement scaling is currently debated [27, 42, 30, 57, 58, 59, 60, 61]. Arguments based on symmetry considerations suggest that such logarithmic phase merely appears as a transient, eventually saturating at a finite area-law value for the entanglement in the thermodynamic limit $L\to\infty$ [58]. If this is indeed the case, our numerical observations would reveal a sharp crossover in the entanglement behavior, as a function of the unraveling, for the analyzed system sizes.
@@
In Ref. [32] an unraveling transition was observed in a fully random quantum circuit, where one of the two “extremal” unravelings could be absorbed in the random Hamiltonian ensemble. In contrast, our system features a deterministic Hamiltonian evolution and two unravelings both substantially altering the unitary dynamics. As a result, the entanglement phase transition we observe cannot be directly mapped onto a “standard” measurement-induced one [32]. It would be interesting to explore whether similar phenomenology might arise in random circuits for quantum communication (see, e.g., Ref. [62]), for example considering different unravelings of the same quantum channel.
Signatures of our unraveling-induced entanglement phase transition may be observed in experiments with cold-atoms [47], for instance by investigating the behavior of many-body correlations [27]. Experimentally, different unravelings may be obtained by varying the quadrature component of the monitored output field [47]. This possibility may allow for controlling entanglement in many-body states via continuous monitoring, which is potentially relevant for tailoring quantum correlations in noisy devices.